OpenStudy (anonymous):

Write an equation in slope-intercept form of the line through point P(8, 2) with slope 4.

OpenStudy (demonchild99):

y=mx +b

OpenStudy (demonchild99):

do you know how to apply that?

OpenStudy (anonymous):

first plug the points and slope into y=mx+b 2=4(8)+b now solve for b b = -30 now plug in the yintercept and the slope y=4x-30

OpenStudy (demonchild99):

quick question.....did that help any?....lol